OpenStudy (jamaljunior):

The equation of a line is −2x+5y=−20−2x+5y=−20. What is the x-intercept of the line?

OpenStudy (kenrick143):

Can you simplify?

OpenStudy (jamaljunior):

idk

OpenStudy (kenrick143):

Why does it have 2 equal sign?

OpenStudy (mathmale):

Your use of the = sign twice in the same expression is problematic. Please copy down and share this problem in simplest form, using ONLY ONE = sign.

OpenStudy (jamaljunior):

sorry the copy paste messed up https://gyazo.com/9441b832203ae74cc1bcae4dc2ee1a7a

OpenStudy (kenrick143):

Letting y=0

OpenStudy (kenrick143):

You can solve the x intercept the same as y intercept letting x=0

OpenStudy (jamaljunior):

how

OpenStudy (kenrick143):

-2x+5(0)=-20

OpenStudy (kenrick143):

Letting y=0 you can now solve for x. :)

OpenStudy (kenrick143):

Got it?

OpenStudy (jamaljunior):

how do i solve for x though

OpenStudy (kenrick143):

-2x+5(0)=-20 ---> 5*0=0 -2x+0=-20 -2x=-20 ---> devide by -2 both sides -1/2(-2x)=-1/2(-20) x=10

OpenStudy (jamaljunior):

thank you.